Carmel Company has a frequent buyer program for its customers, where the customers can attain an “elite” level based on the number of orders and the total revenue of the orders. There are two elite levels: Platinum and Titanium. The benefits of elite membership include discounts and access to special customer service representatives who can resolve problems. The company has one full-time customer representative per 200 Titanium customers and one full-time customer representative per 2,000 Platinum customers. Customer representatives receive salaries plus bonuses of 2 percent of customer gross margin. Carmel spends 80 percent of its promotion costs on Titanium customers to encourage their loyalty.

  

Customer Costs Total Titanium Platinum
Number of customers 27,000 7,000 20,000
Average customer representative salary $ 60,000 $ 60,000
Promotion costs $ 2,050,000
Average gross margin per customer $ 1,510 $ 285

   

Required:

a. Calculate the totals of the items below for both titanium and platinum customers, as well as the excess of gross margin over customer costs for each category.

b. Which customers are more profitable?
Answer #1

SOLUTION

A.

Titanium Platinum
Total gross margin (a) (7,000*$1,510) , (20,000*$285) 10,570,000 5,700,000
Customer representative salaries & bonuses (b) 2,311,400 714,000
Promotion costs (c) (2,050,000*80%), (2,050,000*20%) 1,640,000 410,000
Excess of gross margin over customer cost (a-b-c) 6,618,600 4,576,000

Working note-

Titanium - ($60,000*7,000/ 200) + (2%*10,570,000)

= 2,100,000 + 211,400 = 2,311,400

Platinum - (60,000*20,000/2,000) + (2%*5,700,000)

= 600,000 + 114,000 = 714,000

B.  Titanium customers are more profitable.